A rocket moves upward starting from rest with an a
Answer #1

a)

Hmax = distance trvalled during constant acceleration and distance trvelled after engine is off

Hmax = 0.5 x20x5x5 + (20x20x5x5/2g)

Hmax = 760.2 m

b)

at maximum height velocity = 0

therefore we have

V^2 -u^2 = 2x g x Hmax

V = 122.065m/s
Answer #2

SOLUTION :


a.



Velocity, v(5)  after 5 sec. = a t = 20*5 = 100 m/s

Average velocity during this interval = (0 + 100)/2 = 50 m/s


As the acceleration during this interval is constant, average velocity can be considered as constant velocity during the period.


So, height H(5) attained in 5 sec. 

= average velocity * time

= 50*5

= 250 m 


After 5 sec. engine is turned off and gravity comes into play.


So, 


Initial velocity = v(5) = 100m/s.


Gravitational acceleration (opposite to the up movement) = - g = - 9.8 m/s^2


At he maximum height, Hm , velocity is 0 m/s.


Let it takes time t sec. to reach the highest point.


So,


v(t) = v(5) - g t 

=> 0 = 100 - 9.8 t 

=> t = 100/9.8 = 10.2041 s


Average velocity in the in the duration of t sec = (100+0)/2 = 50 m/s


Height attained in t sec, H(t) 

= average velocity * t 

= 50 t 

= 50 * 10.2041

= 510.20 m


So,


Maximum height, Hm

= H(5) + H(t) 

= 250 + 510.20

= 760.20 m (ANSWER).


b.


From maximum height it starts falling. 


Initial velocity = 0 m/s


Gravitational acceleration = g = 9.8 m/s^2


Distance of fall = 760.20 m


Let it take t’ sec to fall to ground and v(t’) be the velocity I t’ sec.


So,


v(t’) = 0 + g t’ = 0 + 9.8 t’ = 9.8 t’ 


Average velocity = (0 + v(t’))/2 = (0 + 9.8 t’)/2 = 4.9 t’


Distance travelled in t’ sec = Avg. velocity * t’ 

=> 760.20 = 4.9 t’ * t’ = 4.9 t’^2

=> t’ = sqrt(760.20/4.9) = 12.4556 s


So, 


Velocity while hitting the ground, v(t’)

= 9.8 * 12.4556

= 122.06 m/s (ANSWER)

answered by: Tulsiram Garg